OpenStudy (anonymous):

open parentheses x to the 5 eighths power close parentheses to the 2 thirds power.

OpenStudy (anonymous):

\[(x^a)^b=x^{ab}\] \[(x^{\frac{5}{8}})^{\frac{2}{3}}=x^{\frac{5}{8}*\frac{2}{3}}\]

OpenStudy (anonymous):

simplify